# Scalable Hypergraph Structure Learning with Diverse Smoothness Priors

### v1

@@ Abstract
In graph signal processing, learning the weighted connections between nodes from a set of sample signals is a fundamental task when the underlying relationships are not known a priori. This task is typically addressed by finding a graph Laplacian on which the observed signals are smooth. With the extension of graphs to hypergraphs – where edges can connect more than two nodes – graph learning methods have similarly been generalized to hypergraphs. However, the absence of a unified framework for calculating total variation has led to divergent definitions of smoothness and, consequently, differing approaches to hyperedge recovery. We confront this challenge through generalization of several previously proposed hypergraph total variations, subsequently allowing ease of substitution into a vector based optimization. To this end, we propose a novel hypergraph learning method that recovers a hypergraph topology from time-series signals based on a smoothness prior. Our approach addresses key limitations in prior works, such as hyperedge selection and convergence issues, by formulating the problem as a convex optimization solved via a forward-backward-forward algorithm, ensuring guaranteed convergence. Additionally, we introduce a process that simultaneously limits the span of the hyperedge search and maintains a valid hyperedge selection set. In doing so, our method becomes scalable in increasingly complex network structures. The experimental results demonstrate improved performance, in terms of accuracy, over other state-of-the-art hypergraph inference methods; furthermore, we empirically show our method to be robust to total variation terms, biased towards global smoothness, and scalable to larger hypergraphs.
@@ I Introduction
Hypergraphs are considered as generalized graphs that capture higher order relationships [1]. While graphs encode pairwise relationships between nodes through edges, the higher order nature of hypergraphs extends node relations to allow an arbitrary number of nodes to be connected by a hyperedge. Figure 1 contains a sample hypergraph displaying these higher order connections. Examples of hypergraph models include email networks [2] and brain connectivity networks [3]. In the application of brain connectivity, nodes are regions of the brain specified by lobe and hemisphere, and hyperedges represent frequency bands. Any number of nodes could be connected in a hyperedge if, within the frequency band, there was an appropriate measurement of event-related coherence based on MEG measurements. Hypergraphs can also be used in scenarios where graphs already exist, but the relations could be extended to higher order. Examples of this lie in image processing including compression and 3-D point cloud processes [4].
Figure 1: Example hypergraph capturing higher order relationships. Each black circle is a node and the colored groupings are hyperedges. The hyperedges range from connecting 2 nodes to 4 nodes simultaneously.
@@
Some of the more prevalent and successful learning methods on traditional graphs are based on maximizing the smoothness of a set of sample signals across the graph structure where smoothness is the property describing the way a signal travels between connected nodes. At the core of recovering the topology is the concept of total variation – the measurement of this smoothness property or, more precisely, a measure of how not smooth a particular signal is where maximizing smoothness is synonymous with minimizing the signal’s total variation. To this end, Dong et al. [8] use a factor analysis model with a Gaussian probabilistic prior to introduce a method of recovering a graph Laplacian matrix that is effective on smooth graph signals.
With inspiration from Dong et al., Kalofolias [9] developed a graph learning method based on smooth signals that infers a graph weighted adjacency matrix. Both of these methods revolve around maximizing graph signal smoothness, or equivalently, minimizing graph total variation. These works are the basis for several other graph learning related literature [10, 11, 12]. Furthermore, the smoothness prior is effective and abundantly used in several other graph learning contexts [13, 14, 15, 16]. Reiterating that hypergraphs can be viewed as generalized graphs, it is natural that establishing a smoothness prior on hypergraph signals would be beneficial for hypergraph structure inference.
Hypergraph learning is an emerging research topic of increasing interest. Some existing works employ the use of hypergraph structure inference, oftentimes for downstream tasks. Gao et al. [17] developed a hypergraph convolution based neural network for representation learning known as HGNN+. This method uses a two step learning process where the first is modeling the hypergraph structure and the second is hypergraph convolution, which can be thought of as part of the downstream task. Gao et al. offers several hyperedge group generation options, but one in particular makes use of node features. In the node feature space, each node’s K-nearest neighbors (K-NN) are selected and formed into a hyperedge with the original node. Alternatively, the K-NN selection can be based on some maximum feature distance.
@@
Zhang et al. [20] propose a dynamic hypergraph learning method that alternates between updating the hypergraph incidence matrix and a label projection matrix. The problem formulation makes use of a matrix based smoothness prior on both the label projections and the node features. Another smoothness based method from Tang et al. [21] performs community detection on a line graph to form hyperedges. They use weighted Euclidean distance as total variation on a graph before proceeding to line graph construction and subsequent hypergraph structure recovery. While the methods of Zhang et al. and Tang et al. identify node signal smoothness to be of importance, the lack of uniformity of the smoothness measurement discourages judgments between which one is more effective, or if it is application dependent.
@@
Now as with graphs, hypergraph smoothness and total variation can be referenced interchangeably with total variation referenced as the smoothness measurement. However, there is disagreement within the hypergraph signal processing literature regarding the exact form of the total variation cost function. This diversity of smoothness measures can be attributed to the lack of a unified hypergraph signal processing framework. Consequently, there has been little work done to investigate the effects of different hypergraph total variations in structure recovery. In doing such research, this would provide insight into the validity of proposed hypergraph smoothness definitions. Case in point, Tang et al. [22] propose the hypergraph learning method Hypergraph Structure Inference Under Smoothness Prior (HGSI). The core of HGSI is a minimization problem that generates a vector-based solution capturing the hypergraph structure. Each term in the objective function controls a different aspect of the resultant hypergraph structure. Prominently discussed is the use of a smoothness prior through a total variation term limited to the metric equivalent of a sum of squared maximum differences between pairwise nodes derived by expansion of the hyperedges.
@@
Following their outline, we propose a novel inference method that updates and improves upon the approach of HGSI in several ways. We allow our smoothness term to accomodate an assortment of total variation definitions taking into account squared and absolute differences between nodes as well as choosing the maximum difference within the set of nodes forming a hyperedge. We also update the degree positivity term to operate on a dual variable. In doing so, the closed form optimal solution of HGSI is avoided and replaced with a primal-dual algorithmic approach, which benefits in several ways including more equitable hyperedge selection. To further control the structural properties of the hypergraph, we also introduce scalar parameters to enhance or diminish the effects of certain terms such as the degree of nodes. Finally, through an investigation of the use of K-NN to reduce the number of hyperedge possibilities, we propose a method that is scalable without compromising the span of the hyperedge search. Subsequently, we avoid limiting the resulting learned structure which followed assumptions that did not always maintain a desirable set of hyperedges. Experimentally, we show our method, across the majority of scenarios and metrics, outperforms HGSI, the K-NN approach of Gao et al., and a tensor based method from Pena-Pena [23] known as PDL-HGSP.
Our work addresses limitations of prior research with the following contributions. First, we formulate a hypergraph learning model based on a smoothness prior. This model is derived from already existing graph and hypergraph learning works, however the final result is novel. We then structure the model such that a variety of smoothness, or total variation, terms can be substituted. In doing so, our work has a broader range of applications. For instance, since the optimization is robust to term substitution, our model can be fine tuned to the problem at hand with the most appropriate total variation term dependent on the desired learned structure. Furthermore, this substitution contributes to testing the efficacy of hypergraph smoothness definitions. Finally, we demonstrate why certain assumptions in previous works are undesirable and design our method to specifically avoid these issues as they relate to hyperedge selection and solution convergence. In doing so, our learning method becomes scalable to larger problems and adheres to a more complete view of the span of hyperedge possibilities.
@@ II Preliminaries
A hypergraph can capture higher order relationships beyond just the pairwise. More formally, a hypergraph $\mathcal{H}=(V(\mathcal{H}),E(\mathcal{H}))$ is made up of a set of nodes $V(\mathcal{H})=\{v_{1},v_{2},\dots,v_{N}\}$ and set of hyperedges $E(\mathcal{H})=\{\mathbf{e}_{1},\mathbf{e}_{2},\dots,\mathbf{e}_{Y}\}$ where a hyperedge $\mathbf{e}_{y}$ is a set that is allowed to connect any number of nodes simultaneously. The number of nodes is represented as $N=|V(\mathcal{H})|$, and $Y=|E(\mathcal{H})|$ is the number of hyperedges. The cardinality of a hyperedge is the number of nodes connected in the hyperedge set. The maximum cardinality $M$ of $\mathcal{H}$ is the largest number of nodes contained in any one hyperedge such that $\mathbf{e}_{y}=\{v_{i},v_{j},\dots,v_{l}:|\mathbf{e}_{y}|\leq M\}$ and $M\leq N$. Hyperedges may be assigned weights indicating the level of correlation between the connected nodes. Larger weights indicate higher correlation, and the system of defining correlation is application dependent.
@@
Figure 2: Different hypergraph representations through graph expansions of the original structure in Figure 1. They are (a) clique expansion and (b) star expansion. The expansion in (a) fails to represent hyperedge e4 while (b) accurately captures it. Therefore, in the absence of e4, the clique expansion would remain the same as in (a) while (b) would update accordingly.
@@
such that $D$ is the count of every possible hyperedge a hypergraph with cardinalities $\mathcal{K}$ could have.
@@
where each element of the pairwise distance vector, $\mathbf{z}$, is $z_{d}=||\mathbf{X}_{i,:}-\mathbf{X}_{j,:}||_{2}^{2}$ and $\mathbf{X}_{i,:}^{\top},\mathbf{X}_{j,:}^{\top}\in\mathbb{R}^{P}$ are graph time-series signals on nodes $v_{i}$ and $v_{j}$, respectively, corresponding to row $i$ and $j$ of $\mathbf{X}$. In parallel, hypergraph total variation can be represented in the same way. In this case, $\mathbf{w}\in\mathbb{R}^{D}$ holds the hyperedge weights and $\mathbf{z}=[z_{\mathbf{e}_{1}},z_{\mathbf{e}_{2}},\dots,z_{\mathbf{e}_{d}}]\in\mathbb{R}^{D}$ holds some form of distance measure for each potential hyperedge. Since hyperedges are allowed to connect an arbitrary number of nodes together, the realization of the pairwise distance vector needs to be updated to a generalized distance measure. This notion has inspired several kinds of hyperedge distance definitions. With each new distance calculation stored in $\mathbf{z}$, there is a new hypergraph total variation measure $\Omega(\mathbf{w})$.
@@
where $s(\mathbf{X}_{i,:},\mathbf{X}_{j,:})$ is a kind of measure of difference between node time-series signals. The function, $t(\cdot)$, then combines these measures across all pairings of nodes $v_{i},v_{j}\in\mathbf{e}_{d}$ into the distance measure which is stored in $\mathbf{z}$. This can then be scaled by the hyperedge weight $w_{\mathbf{e}_{d}}$, which is from $\mathbf{w}$. Finally, $T(\cdot)$ combines all measures across all hyperedges from $E(\mathcal{H})$. In its entirety, the unified smoothness term is defined as $\Phi(\mathbf{X})$. We find that the smoothness measurements we use follow the unified form of (4) and therefore are valid hypergraph total variations.
TABLE I: Calculated total variations for each perturbed hypergraph signal of Figure 3.
Total Variation Term No Perturbation Total Variation Small Perturbation Total Variation Large Perturbation Total Variation
Sum-Square 71.13 73.00 114.6
Sum-Absolute 25.36 25.45 31.78
Max-Absolute 12.80 12.80 14.47
Max-Square 44.64 44.64 62.61
A first hypergraph total variation of interest to learning hypergraphs is the Sum-Square variation found in the work from Agarwal et al. [25] defined according to
@@
which uses the squared distances between node time-series. The summations are across all pairs $v_{i}$ and $v_{j}$ in the hyperedge of interest, $\mathbf{e}_{d}$, and then across all hyperedges, $\mathbf{e}_{d}\in E(\mathcal{H})$. When $M=2$, this smoothness measure decomposes to the traditional means of measuring total variation in graphs, derived by either star or clique expansions. Figure 2a and 2b contain an example clique and star expansion, respectively, of the hypergraph of Fig. 1. The clique expansion is considered only as an approximation as it cannot represent non-isomorphic hypergraphs with the same clique expansion [26]. More simply, if two hypergraphs have different structures, the clique expansion could represent them in the exact same way as demonstrated in Fig. 2. The star expansion demonstrates the ability to accurately represent all hyperedges, and as such is considered a better representation.
The second hypergraph total variation of interest from Nguyen and Mamitsuka [24], noted to be associated with the clique expansion, is the Sum-Absolute variation defined as
@@
which is similar to (5) except the difference measure is now the absolute value. Both (5) and (6) sum the differences between node time-series for all $v_{i},v_{j}\in\mathbf{e}_{d}$. This property is not advantageous when combined with a minimization problem such as (1). The main issue is that as the cardinality of a hyperedge increases, there are more pairs of $v_{i},v_{j}$ and, therefore, more summations of time-series differences increasing the total variation. With no sort of normalization or averaging, this causes a magnitude imbalance between the contribution of hyperedges with mixed cardinalites to the overall total variation.
@@
which injects some nonlinearity by taking the maximum absolute difference between node time-series within a hyperedge. The use of the $\max(\cdot)$ difference avoids the magnitude imbalance issue and provides an interesting notion of hypergraph smoothness. This implies that the total variation contribution of a hyperedge is governed by the two nodes with the least similarity, or largest variance.
The use of the absolute difference could provide issues when it comes to a minimization problem as any perturbation to the smoothness calculation could result in changes that are too small to make a significant impact. To counteract this, we introduce the last hypergraph total variation to be the Max-Square variation defined by
@@
Figure 3: Example hypergraph with perturbed signal on node 5. Colorings on node 5 from top to bottom represent: (top) no perturbation, (middle) a small perturbation, and (bottom) a large perturbation.
To highlight the structural properties of each total variation term, we refer to the example of Fig. 3 where the topology of the hypergraph is the same as in Fig. 1, but here, we include a randomly generated hypergraph signal from a multivariate Gaussian distribution with mean 0 and standard deviation 2. That is to say, the signal was not designed to reflect the hyperedge connections, only to demonstrate total variation properties. We introduce changes, or perturbations, to the hypergraph signal, but only on node 5. There are three coloring representations for node 5 which correspond to the original, small perturbed, and large perturbed signals. By only perturbing node 5, we can calculate the total variation using each proposed term and observe the system response. Table I contains a report of the calculated total variation for each resultant hypergraph.
The most appropriate way of analyzing the total variations of Table I is to refer to their relative change from the original hypergraph to the perturbed hypergraphs. From the small signal change between the orginal and small perturbed signals, we see that only the total variation terms based on the summation of differences change while the ones based on a maximum difference remain the same. This is because node 5 is not part of the difference measure that creates the max distance between two nodes in the hyperedge. Furthermore, the small perturbation was not large enough to push node 5 into the max difference. We also note that the change in total variation using the Sum-Absolute was much smaller than the change using Sum-Square, which can be attributed to the use of a squared term versus a piecewise linear term.
Looking at the change between the original and large perturbed signals, we see much larger total variation differences caused by the larger perturbation. This time, the signal change was large enough to push node 5 into the max distance calculation which in turn caused the Max-Square and Max-Absolute total variations to increase. Furthermore, the change in total variation of the terms with square differences are once again much larger than those with absolute differences. Therefore, we can say that total variations with absolute differences will represent a signal to be smoother than those with squared distances. This is an important differentiation because these total variation terms will end up being substituted into a minimization problem. If the total variation term has a squared difference, then the optimization will be more sensitive to small changes in the input signals. However, the Max-Square total variation demonstrated a smaller change than the Sum-Square, which indicates that even among the use of squared differences, the Sum-Square will be the total variation most sensitive to signal perturbations.
III HGSI
Tang et al. [22] proposed the hypergraph structure inference method (HGSI) to make use of the Max-Square variation, through $\Omega(\mathbf{w})=\mathbf{z}^{\top}\mathbf{w}$, and includes the minimization

### v2

@@ Abstract
In graph signal processing, learning the weighted connections between nodes from a set of sample signals is a fundamental task when the underlying relationships are not known a priori. This task is typically addressed by finding a graph Laplacian on which the observed signals are smooth. With the extension of graphs to hypergraphs – where edges can connect more than two nodes – graph learning methods have similarly been generalized to hypergraphs. However, the absence of a unified framework for calculating total variation has led to divergent definitions of smoothness and, consequently, differing approaches to hyperedge recovery. We confront this challenge through generalization of several previously proposed hypergraph total variations, subsequently allowing ease of substitution into a vector based optimization. To this end, we propose a novel hypergraph learning method that recovers a hypergraph topology from time-series signals based on a smoothness prior. Our approach, designated as Hypergraph Structure Learning with Smoothness (HSLS), addresses key limitations in prior works, such as hyperedge selection and convergence issues, by formulating the problem as a convex optimization solved via a forward-backward-forward algorithm, ensuring guaranteed convergence. Additionally, we introduce a process that simultaneously limits the span of the hyperedge search and maintains a valid hyperedge selection set. In doing so, our method becomes scalable in increasingly complex network structures. The experimental results demonstrate improved performance, in terms of accuracy, over other state-of-the-art hypergraph inference methods; furthermore, we empirically show our method to be robust to total variation terms, biased towards global smoothness, and scalable to larger hypergraphs.
@@ I Introduction
Hypergraphs are considered as generalized graphs that capture higher order relationships [1]. While graphs encode pairwise relationships between nodes through edges, the higher order nature of hypergraphs extends node relations to allow an arbitrary number of nodes to be connected by a hyperedge. Figure 1 contains a sample hypergraph displaying these higher order connections where nodes are considered workers and hyperedges connect workers who are collaborating on a project. Hypergraph models have been applied in the contexts of email networks [2], multi-cell multi-user power allocation [3], machinery fault diagnosis [4], short text classification [5], traffic forecasting [6], and brain connectivity networks [7]. In the application of brain connectivity, nodes are regions of the brain specified by lobe and hemisphere, and hyperedges represent frequency bands. Any number of nodes could be connected in a hyperedge if, within the frequency band, there was an appropriate measurement of event-related coherence based on MEG measurements. Hypergraphs can also be used in scenarios where graphs already exist, but the relations could be extended to higher order. Examples of this lie in 3-D point cloud processes [8], object classification [9], pixel segmentation [10], and noise reduction [11].
Figure 1: Example of a hypergraph network capturing higher order relationships. Nodes A through F are workers, and the colored connections e1 through e4 are hyperedges relating workers who are on the same project together.
@@
Some of the more prevalent and successful learning methods on traditional graphs are based on maximizing the smoothness of a set of sample signals across the graph structure where smoothness is the property describing the way a signal travels between connected nodes. At the core of recovering the topology is the concept of total variation (TV) – the measurement of this smoothness property or, more precisely, a measure of how not smooth a particular signal is where maximizing smoothness is synonymous with minimizing the signal’s total variation. To this end, Dong et al. [15] use a factor analysis model with a Gaussian probabilistic prior to introduce a method of recovering a graph Laplacian matrix that is effective on smooth graph signals.
Figure 2: Example of a simplicial complex representing a similar higher order network as in Fig. 1 with nodes being workers. The key difference here is that Fig. 1 has a structure of four hyperedges while this network has a 0-simplex for every node, a 1-simplex for every line (dashed or solid lines), a 2-simplex for every triangle (colored faces), and a 3-simplex for every tetrahedron (groups of 4 nodes). Each hyperedge has been converted into its highest order simplex containing all lower order simplices.
With inspiration from Dong et al., Kalofolias [16] developed a graph learning method based on smooth signals that infers a graph weighted adjacency matrix. Both of these methods revolve around maximizing graph signal smoothness, or equivalently, minimizing graph total variation. These works are the basis for several other graph learning related literature [17, 18, 19]. Furthermore, the smoothness prior is effective and abundantly used in diverse graph learning contexts [20, 21, 22, 23]. Reiterating that hypergraphs can be viewed as generalized graphs, it is natural that establishing a smoothness prior on hypergraph signals would be beneficial for hypergraph structure inference. However, while total variation on graphs have a well established connection to the notion of frequency in the graph domain, with smaller total variation indicating a lower frequency signal, the spectral relationship between hypergraphs and frequency via total variation is still an area of emerging research [24, 25, 26, 27].
Some existing hypergraph learning works employ the use of hypergraph structure inference for downstream tasks. Gao et al. [28] developed a hypergraph convolution based neural network for representation learning known as HGNN+. This method uses a two step learning process where the first is modeling the hypergraph structure and the second is hypergraph convolution, which can be thought of as part of the downstream task. Gao et al. offers several hyperedge group generation options, but one in particular makes use of node features. In the node feature space, each node’s K-nearest neighbors (K-NN) are selected and formed into a hyperedge with the original node. Alternatively, the K-NN selection can be based on some maximum feature distance. In a separate example, Yang et al. [29] implement the Volterra model to estimate higher-order connections through use of structural equation models of nonlinear relationships. This learned structure is subsequently used in the task of link prediction for social networks.
@@
Zhang et al. [32] propose dynamic hypergraph learning that alternates between updating the hypergraph incidence matrix and a label projection matrix. The problem formulation makes use of a matrix based smoothness prior on both the label projections and the node features. Another smoothness based method from Tang et al. [33] performs community detection on a line graph to form hyperedges. They use weighted squared Euclidean distance as total variation on a graph before proceeding to line graph construction and subsequent hypergraph structure recovery. While the methods of Zhang et al. and Tang et al. identify node signal smoothness to be of importance, the lack of uniformity of the smoothness measurement discourages judgments between which one is more effective, or if it is application dependent.
Simplicial complexes are hierarchical structures that are an alternative for higher order network representation. Similar to hypergraphs, simplicial complexes can group together any number of nodes in a $c$-simplex where $c$ is one less than the number of nodes in the simplex, representing its geometric dimension. However, simplicial complexes diverge from hypergraph structures in that any $c$-simplex must also include all subset simplices from dimension $c-1$ to 00 [34]. This inherently limits the representational ability of simplicial complexes to densely connected spaces where all $c$ dimensions must be considered. Hypergraphs do not share this property, and are free to group any number of nodes in any subset without limitation. Figures 1 and 2 illustrate this comparison where the simplicial complex represents the same highest order connections as the hypergraph, but must also include all unique subsets in the topology. Furthermore, a hypergraph could represent a simplicial complex, and therefore is the more robust choice of network structure, at the sacrifice of scalability. So, while simplicial complexes are valid higher order structures that have been inferred from nodal signals [35, 36, 37], this work will focus on hypergraph recovery since hypergraphs are inherently different structures [34] which have a broader range of applications.
@@
Now as with graphs, hypergraph smoothness and total variation can be referenced interchangeably with total variation being the measurement of smoothness. However, there is disagreement within the hypergraph signal processing literature regarding the exact form of the total variation cost function. This diversity of smoothness measures can be attributed to the lack of a unified hypergraph signal processing framework. Consequently, there has been little work done to investigate the effects of different hypergraph total variations in structure recovery. In doing such research, this would provide insight into the validity of proposed hypergraph smoothness definitions. Case in point, Tang et al. [38] propose the hypergraph learning method Hypergraph Structure Inference Under Smoothness Prior (HGSI). The core of HGSI is a minimization problem that generates a vector-based solution capturing the hyperedges. Each term in the objective function controls a different aspect of the resultant hypergraph structure. Prominently discussed is the use of a smoothness prior through a total variation term limited to the metric equivalent of a sum of squared maximum differences between pairwise nodes derived by expansion of the hyperedges.
@@
Following their outline, we propose a novel inference method known as Hypergraph Structure Learning with Smoothness (HSLS). Our method updates and improves upon the approach of HGSI in several ways. We allow our smoothness term to accommodate an assortment of total variation definitions taking into account linear summations and nonlinear maximums of squared and absolute differences between nodes. We also update the degree positivity term to operate on a dual variable. In doing so, the closed form optimal solution of HGSI is avoided and replaced with a primal-dual algorithmic approach, which benefits in several ways including more equitable hyperedge selection. Scalar parameters are introduced to enhance or diminish the effects of certain terms such as the degree of nodes. Finally, through an investigation of the use of K-NN to reduce the number of hyperedge possibilities, we propose a method that is scalable without compromising the span of the hyperedge search. Subsequently, we avoid limiting the resulting learned structure which followed assumptions that did not always maintain a desirable set of hyperedges. Experimentally, we show our method, across the majority of scenarios and metrics, outperforms HGSI, the K-NN approach of Gao et al., and a tensor based method from Pena-Pena [26] known as PDL-HGSP. We also demonstrate our method’s ability to prioritize smooth structure recovery and improve upon a pre-existing learning application.
Our work addresses limitations of prior research with the following novel contributions:
  1. 1.
    We formulate a hypergraph learning model HSLS based on a smoothness prior which improves upon existing hypergraph learning works.
  2. 2.
    HSLS is structured such that a variety of smoothness, or total variation, terms can be substituted, such that we can test the efficacy of hypergraph smoothness definitions in varying applications.
  3. 3.
    We demonstrate why certain assumptions in previous works are undesirable and design HSLS to specifically avoid these issues as they relate to hyperedge selection and solution convergence. In doing so, our learning method becomes scalable to larger problems and adheres to a more complete view of the span of hyperedge possibilities.
  4. 4.
    We generate and analyze original experimental results using HSLS in three application domains, each providing new empirical results both for the proposed method and smoothness definitions.
The rest of the paper is organized as follows: Section II covers preliminary information including hypergraph notation, optimization, and total variation measures. Section III reviews a previous work of hypergraph structure inference and includes a discussion of certain shortcomings their method contains. Then, Section IV introduces the updated novel method HSLS which comprises of a cost function, solved via a convex optimization algorithm, and a hyperedge reduction technique. Section V experimentally evaluates the ability of HSLS to recover accurate and reasonable hypergraph structures and maintain the scalability of the method. Section VI summarizes our findings, and the Appendix contains proofs and derivations.
@@ II Preliminaries
A hypergraph can capture higher order relationships beyond just the pairwise. More formally, a hypergraph $\mathcal{H}=(V(\mathcal{H}),E(\mathcal{H}))$ is made up of a set of nodes $V(\mathcal{H})=\{v_{1},v_{2},\dots,v_{N}\}$ and set of hyperedges $E(\mathcal{H})=\{\mathbf{e}_{1},\mathbf{e}_{2},\dots,\mathbf{e}_{Y}\}$ where a hyperedge $\mathbf{e}_{y}$ is a set that is allowed to connect any number of nodes simultaneously. The number of nodes is represented as $N=|V(\mathcal{H})|$, and $Y=|E(\mathcal{H})|$ is the number of hyperedges. The cardinality of a hyperedge is the number of nodes connected in the hyperedge set. The maximum cardinality $M$ of $\mathcal{H}$ is the largest number of nodes contained in any one hyperedge such that $\mathbf{e}_{y}=\{v_{i},v_{j},\dots,v_{l}:|\mathbf{e}_{y}|\leq M\}$ and $M\leq N$. Hyperedges may be assigned weights indicating the strength of relation between the connected nodes. Larger weights indicate stronger relations, and the system of defining these relationships is application dependent.
@@
Figure 3: Different hypergraph representations through graph expansions of the original structure in Figure 1. They are (a) clique expansion and (b) star expansion. The expansion in (a) fails to distinguish between nodes F and G in hyperedge e3 and e4 as the clique expansion aggregates this information into one representative edge (the red and purple line). Expansion (b) accurately captures this difference. Therefore, in the absence of e4, the clique expansion would remain structurally the same as in (a) while (b) would update accordingly.
@@
such that $D$ is the count of every possible hyperedge a hypergraph with cardinalities $\mathcal{K}$ could have and $L=|\mathcal{K}|$.
@@
where each element of the pairwise distance vector, $\mathbf{z}$, is $z_{d}=||\mathbf{X}_{i,:}-\mathbf{X}_{j,:}||_{2}^{2}$ and $\mathbf{X}_{i,:}^{\top},\mathbf{X}_{j,:}^{\top}\in\mathbb{R}^{P}$ are graph time-series signals on nodes $v_{i}$ and $v_{j}$, respectively, corresponding to row $i$ and $j$ of $\mathbf{X}$. In parallel, hypergraph total variation can be represented in the same way. In this case, $\mathbf{w}\in\mathbb{R}^{D}$ holds the hyperedge weights and $\mathbf{z}=[z_{\mathbf{e}_{1}},z_{\mathbf{e}_{2}},\dots,z_{\mathbf{e}_{D}}]\in\mathbb{R}^{D}$ holds some form of distance measure for each potential hyperedge. Since hyperedges are allowed to connect an arbitrary number of nodes together, the realization of the pairwise distance vector needs to be updated to a generalized distance measure. This notion has inspired several kinds of hyperedge distance definitions. With each new calculation stored in $\mathbf{z}$, there is a new hypergraph total variation measure $\Omega(\mathbf{w})$.
@@
where $s(\mathbf{X}_{i,:},\mathbf{X}_{j,:})$ is a kind of measure of difference between node time-series signals. The function, $t(\cdot)$, then combines these measures across all pairings of nodes $v_{i},v_{j}\in\mathbf{e}_{d}$ into the distance measure which is stored in $\mathbf{z}$. This can then be scaled by the hyperedge weight $w_{\mathbf{e}_{d}}$. Finally, $T(\cdot)$ combines all measures across all hyperedges from $E(\mathcal{H})$. In its entirety, the unified smoothness term is defined as $\Phi(\mathbf{X})$. We find that the smoothness measurements we use follow the unified form of (4) and therefore are valid hypergraph total variations.
It is worth recognizing that there is a wide selection of total variation definitions on hypergraphs. The list is rather expansive, especially considering generalized frameworks like that of (4) and the between-hyperedge $\ell_{p}$ norm [39] defined as
$\left(\sum_{\mathbf{e}_{d}\in E(\mathcal{H})}\sum_{v_{i},v_{j}\in\mathbf{e}_{d}}||\mathbf{X}_{i,:}-\mathbf{X}_{j,:}||_{p}\right)^{\frac{1}{p}},$ (5)
where (5) is similar to the p-regularizer used in undirected graph works [40, 41]. There are many combinations of functions that can be substituted into (4) and many values of $p$ that can be used in (5). To this end, we have chosen a group of valid total variation terms that are able to be cast in the form of (3). The chosen smoothness terms are by no means an exhaustive list; however, they do cover a diverse range of linear and nonlinear forms and common $\ell_{p}$ norms. Furthermore, all definitions fit (4), and some also satisfy (5).
We now introduce the four total variation measures used in this work, all of which are convex operations. A first hypergraph total variation of interest to learning hypergraphs is the Sum-Square variation found in the work from Agarwal et al. [42] defined according to
@@
which uses the squared Euclidean distances between node time-series. The summations are across all pairs $v_{i}$ and $v_{j}$ in the hyperedge of interest, $\mathbf{e}_{d}$, and then across all hyperedges, $\mathbf{e}_{d}\in E(\mathcal{H})$. When $M=2$, this smoothness measure decomposes to the traditional means of measuring total variation in graphs, and can be derived by either star or clique expansions. Figure 3a and 3b contain an example clique and star expansion, respectively, of the hypergraph of Fig. 1. The clique expansion is considered only as an approximation as it cannot represent non-isomorphic hypergraphs with the same clique expansion [43]. More concretely, if there are two hyperedges where one has at least two nodes that are the subset of the other, then the clique expansion would have at least one edge with a multiplicity of 2. It is standard for the clique expansion to aggregate edges with multiplicity greater than 1 into a single representative edge [43, 44, 45], in which case the distinction between hyperedges is lost. Alternatively, the star expansion demonstrates the ability to accurately represent all hyperedges, and as such is considered a better representation. This concept is demonstrated in the example of Fig. 3.
The second hypergraph total variation of interest from Nguyen and Mamitsuka [39], noted to be associated with the clique expansion and align with the framework of (5) when $p=1$, is the Sum-Absolute variation defined as
@@
which is similar to (6) except the difference measure is now the absolute value. Both (6) and (7) sum the differences between node time-series for all $v_{i},v_{j}\in\mathbf{e}_{d}$. This property is not advantageous when combined with a minimization problem such as (1). The main issue is that as the cardinality of a hyperedge increases, there are more pairs of $v_{i},v_{j}$ and, therefore, more summations of time-series differences increasing the total variation. With no sort of normalization or averaging, this causes a magnitude imbalance between the contribution of hyperedges with mixed cardinalities to the overall total variation.
@@
which injects some nonlinearity by taking the maximum absolute difference between node time-series within a hyperedge. This total variation also follows the framework of (5) where the inner summation is allowed to be replaced with the nonlinear maximum operation, and $p=1$. The use of the max difference avoids the magnitude imbalance issue and provides an interesting notion of hypergraph smoothness. This implies that the total variation contribution of a hyperedge is governed by the two nodes with the least similarity, or largest variance.
The use of the absolute difference in a minimization problem may be problematic as any perturbation to the smoothness calculation could result in changes that are too small to have a meaningful effect. To counteract this, we introduce the last hypergraph total variation to be the Max-Square variation defined by
@@
III Hypergraph Structure Inference Under Smoothness Prior (HGSI)
Tang et al. [38] proposed the hypergraph structure inference method HGSI to make use of the Max-Square variation, through $\Omega(\mathbf{w})=\mathbf{z}^{\top}\mathbf{w}$, and includes the minimization